Suppose that the probability that a particular brand of light bulb fails before 900 hours of use is 0.2. If you purchase 3 of these light bulbs, what is the probability that at least one of them lasts 900 hours or more?
Answer by stanbon(75887)

Suppose that the probability that a particular brand of light bulb fails before 900 hours of use is 0.2. If you purchase 3 of these light bulbs, what is the probability that at least one of them lasts 900 hours or more?
------
P(at least one in 3) = 1 - P(none of 3 lasts 900 hrs)
---
= 1 - (0.2)^3
---
= 1 - 0.008
---
= 0.992
=========
Cheers,
Stan H.